Tan= square root(3)/3 find sec
Answer by ewatrrr(24785)

Tan (π/6) = √3/3 and Tan (7π/6) = √3/3 Interval 0 to 2π
sec(π/6) = 1/cos(π/6) = 1/(√3/2) = 2/√3 0r 2√3/3
sec(7π/6) = 1/cos(7π/6) = 1/-(√3/2) = -2/√3 0r -2√3/3
below (cosx, sinx) Summary